OpenStudy (anonymous):

solve for the variable -22 = x - 7 I solved and got 29 but it was incorrect, can anyone correct me?

OpenStudy (lgbasallote):

i can

OpenStudy (lgbasallote):

-22 = x - 7 add + 7 to both sides -22 + 7 = x - 7 + 7 -22 + 7 = x is this what you did?